The Van Der Waals constants 'a' and 'b' for NH₃ (ɡ) are 0.4225 Pa m⁶ mol⁻² and 37.07×10⁶ mol⁻¹ respectively . calculate the values of its critical constants .
1
Expert's answer
2019-10-28T07:49:24-0400

Citical temperature "T_c=\frac{8a}{27Rb}=0.4\times10^{-9}"

Critical volume "V_c=3b=1.1121\times10^8"

Critical pressure "P_c=\frac{a}{27b^2}=0.113\times10^{-16}"